Solve for each system by the substitution method
6x - 3y = 3
10x = y + 7
..
10x=[y]+7
-7 -7
----------
10x-7=y
+7 +7
-----------
10=y-7
+7 +7
y=17
Plug it in the first equation
6x-3(17)=y
6x-51=3
+51 +51
--------
6x=54
/6
x=9
Check:
6(9)-3(17)=3
54-51=3
3=3

Answer by Bebita1992(4)

Solve for each system by the substitution method
6x - 3y = 3
10x = y + 7
..
10x=[y]+7
-7 -7
----------
10x-7=y
+7 +7
-----------
10=y-7
+7 +7
y=17
Plug it in the first equation
6x-3(17)=y
6x-51=3
+51 +51
--------
6x=54
/6
x=9
Check:
6(9)-3(17)=3
54-51=3
3=3